OpenStudy (anonymous):

Suppose the spring in this spring-powered rifle has a spring constant of k = 1,000 N/m. How much work can it do in accelerating the projectile, if it is compressed by x = 0.10 m? 20 J 5.0 × 103 J 10 J 5.0 J

OpenStudy (anonymous):

this picture goes with the question

OpenStudy (anoop27):

w=1/2Kx^2 =0.5*1000*0.1*0.1

OpenStudy (anonymous):

so the answer is D. 5.0 J

OpenStudy (anonymous):

right?

OpenStudy (anoop27):

yes

OpenStudy (anoop27):

@megen505 yes ans is D

OpenStudy (anonymous):

perfect thank you so much!

OpenStudy (anoop27):

ur welcome